A college charters a bus for $1300 to take a group to a museum. When fourteen more students join the trip, the cost per student drops by $17.50. How many students were in the original group?
Answer by stanbon(75887)

A college charters a bus for $1300 to take a group to a museum. When fourteen more students join the trip, the cost per student drops by $17.50. How many students were in the original group?
-------------------------------------------
Let the original # in the group be "x".
Cost per students was 1300/x dollars.
--------------------------------------------
The # in the revised group is (x+14).
Cost per student is 1300/(x+14) dollars.
--------------------------------------------
Equation:
old average - new average = 17.50
-----------------------------
1300/x - 1300/(x+14) = 17.50 dollars
-------
Multiply thru by x(x+14) to get:
1300(x+14) - 1300x = 17.5x(x+14)
1300*14 = 17.5x^2 + 245x
17.5x^2 + 245x - 18200 = 0
17.5(x^2 + 14x - 1040) = 0
(x+40)(x-26) = 0
---
Positive solution:
x = 26 (# in the original group)
=====================================
Cheers,
Stan H.